What is the solution for:
3x+5y=10
x-5y=-10

Answer by ewatrrr(24785)

 
Hi,
solving by Elimination
3x+5y=10
x-5y=-10
4x = 0
x = 0 and y =2 (0-5y=-10 and -5y=-10, y = 2 )
CHECKING our Answer***Using 3x+5y=10
3·0 + 5·2 = 10